b) When a 4-sided die is rolled, let A be the event that a 1 is rolled and B be the event that an odd number is rolled. i Without doing any calculations, explain why P(A) s P(B) ii) Do the calculations.

Thank you!
Answer #1

(i)

Sample space of A is = 1

Sample space of B is = 1, 3 and 5

So, sample space of A is subset of sample space of B, which means,

ACB

This implies, P(A) <= P(B)

(ii)

Probability of 1 is rolled, P(A) = 1 / 6

Probability of any odd number is rolled, P(B) = 3 / 6 = 1/2

So, P(B) > P(A)

** If this answer does not match please comment